OpenStudy (anonymous):

How many grams of Cl are in 535 g of CaCl2?

OpenStudy (chmvijay):

can u tell me molar mass of CaCl2 and atomic mass of Cl

OpenStudy (anonymous):

The molar mass is 110.98 and the atomic mass of Cl is 35.45

OpenStudy (chmvijay):

110.98 gram of Cacl2 contains =35.45 *2 gram of Cl 535 gramof CaCl2 =35.45*2*535 /110.98=

OpenStudy (anonymous):

ok thank you so much!

OpenStudy (chmvijay):

WELCOME TO OPEN STUDY :)